# Active Learning Methods for Efficient Data Utilization and Model Performance Enhancement thanks:

### v1

@@ IV Evaluations
Evaluation Metrics and Benchmarks for Active Learning: there have been calls from recent research for sound and reproducible evaluation, with appropriate re-benchmarking of available benchmarks in order to yield more sound results [26]. New benchmarks like ALdataset [73], OpenAL [13], and CDALBench [74] facilitate end-to-end comparisons.
A central theme is the recommendation of practical metrics for restoring trust in AL among such studies [4]. Metrics like interpretability metrics and statistical analysis techniques provide greater insight into AL performance. Cross-domain validation and large-scale reproducibility guarantee that methods generalize from one dataset [74]. Furthermore, interpretability and transparency are essential. Studies demand metrics that explain model behavior [13, 4]. Open experimentation environments and shared code enhance reproducibility [26, 73].
Perspectives on metrics and settings differ. Some researchers aim to design new benchmarks and evaluation metrics [73], e.g., OpenAL, proposed by Jonas et al. [13]. Others advocate using existing metrics with new analysis techniques [4]. Experiment settings also vary between studies, from differences in the choice of datasets, oracle implementation, and repetition count [74, 26]. Methodologies vary from benchmark design to statistical analysis. The design of new benchmarks like ALdataset and OpenAL necessitates attention to experimental settings and evaluation metrics [73, 13]. Open-sourced implementations as well as efficient oracle implementations are also important in enabling the comparison of AL approaches [26, 74].
In spite of the advancements made in the construction of evaluation metrics and benchmarks for AL, there are still limitations and challenges. A deficiency in the standardization of benchmarks and evaluation metrics prevents comparison of AL approaches, and variability in the experimental conditions can give rise to conflicting results [26, 73, 75, 76]. The necessity of performing large-scale evaluation with repeated runs highlights the importance of making thorough and trustworthy assessments of AL approaches [74].
In summary, working towards evaluation metrics and benchmarks for AL remains an important research focus, with themes such as the necessity for trustworthy benchmarks, interpretability, cross-domain evaluation, and reproducibility. Metrics and setups for experiments vary among researchers. The technical approaches from these works provide a strong base for subsequent research. Overcoming the limitations and challenges of these works will be vital in furthering our knowledge of AL strategies and their applications in other domains [77, 78].
@@ V Applications
Applications of Active Learning in Computer Vision: AL is becoming more popular in computer vision because it helps reduce the cost of data labeling [16]. Many recent studies focus on lowering annotation costs while keeping training efficient [79, 60]. For example, learning how confident a user is when labeling can help avoid asking for labels that are not necessary, which improves the overall process [79]. Some methods first use weak labels and only request strong labels later. This helps save effort in tasks like object detection [60].
@@
Active learning has many applications in computer vision, e.g. image classification, object detection, and semantic segmentation [80, 58]. New methods aim to make active learning more effective and faster, including uncertainty sampling, query-by-committee, and reinforcement learning [81, 82]. Several papers also discuss how to evaluate active learning better. They suggest new evaluation metrics and testing protocols for realistic comparisons [1, 41].
@@
Take Greedy Active Learning (GAL) for example. Bar et al. [83] propose a greedy-based active learning framework that operates in two stages during each AL cycle. Denote the classifier at iteration $t$ as $\mathcal{C}_{t}$. In the initial stage, a candidate subset $X_{c}\subseteq X_{u}$ of size $K:=|X_{c}|$ is selected from the unlabeled data pool. This subset may either encompass the entire unlabeled set or be filtered based on the top-K relevance scores. Due to inherent class imbalance, $X_{c}$ typically contains many irrelevant samples. In the following stage, a batch $X_{b}\subset X_{c}$ is chosen using an AL strategy. A human oracle then labels the instances in $X_{b}$, updating the labeled set $(X_{l},\mathcal{Y}_{l})$ with their features and labels. A new classifier $\mathcal{C}_{t+1}$ is trained on this augmented dataset for the next iteration.
@@
Researchers are questioning traditional active learning methods. Some propose new frameworks that combine semi-supervised learning and data augmentation [41]. Others explore combining active learning with transfer learning and self-supervised learning [84, 85]. Deep learning models like CNNs and visual transformers are also widely used in active learning for computer vision tasks [86, 87].
@@
Despite recent progress, some challenges still exist. It is hard to scale active learning to very large datasets or complex problems [88, 83]. Also, poor-quality data or labels can hurt performance. Researchers are working on more efficient algorithms to solve these issues [89]. High-quality datasets and better annotation protocols are also important for success [3].
@@
In conclusion, AL plays an important role in computer vision. It helps reduce labeling costs and improve learning efficiency. Recent work explores new evaluation methods and combines active learning with other learning techniques. Researchers continue to improve active learning approaches [90, 91] to mitigate the challenges. The field is growing, and active learning will likely become even more important in future computer vision systems.
Algorithm 3 Greedy Active Learning (GAL) Algorithm [83]
@@
         $x^{*},\hat{l}^{*}\leftarrow\textsc{Next}(\mathcal{X}_{c},\mathcal{X}_{l},\mathcal{Y}_{l})$triangleright Find the point that maximizes impact $\mathcal{S}$
@@
         $\mathcal{S}_{i},\hat{l}_{i}\leftarrow\psi(x_{i},\mathcal{X}_{l},\mathcal{Y}_{l})$triangleright Acquisition function (see Alg. 4)
@@
     return $x_{i^{*}},\hat{l}_{i^{*}}$triangleright Return optimal point and pseudo label
@@
Algorithm 4 Acquisition Functions [83]
function $\psi_{svm}$($x_{i}$, $\mathcal{X}_{l}$, $\mathcal{Y}_{l}$$\triangleright$ SVM
@@
function $\psi_{gp}$($x_{i}$, $\mathcal{X}_{l}$, $\mathcal{Y}_{l}$$\triangleright$ Gaussian Process
@@
Active Learning for Natural Language Processing Tasks: many papers show AL can help improve data efficiency and model accuracy [5, 92, 93]. One important idea is adapting pre-trained language models to the specific task. Training more on unlabeled data can improve results and reduce the need for labels [92]. For example, ensemble methods like Majority-CRF [93] are used in natural language understanding. They reduce error rates by 6.6%–9% compared to random sampling. Another method, active query k-means, also helps increase accuracy and cut training costs in text classification [94].
Model adaptation is a crucial topic in NLP active learning. It shows the importance of learning strategies for pre-trained models [5, 92]. Ensemble methods are popular too, as they help pick the most useful samples [93]. Semi-supervised learning is often used with active learning to reduce labeling needs [95, 94]. Beatty et al. show that choosing when to stop active learning using labeled vs. unlabeled data is important, especially for text classification [96]. Good evaluation tools, e.g., datasets and metrics, are also important. Most papers use F1-score to measure performance [95]. The corresponding equations in pseudocodes are:
@@
There are different opinions on which query strategy is best. Some focus on exploitation, some on exploration, and others mix both [95]. Some researchers prefer continuous training on unlabeled data to adapt Language Models (LMs). Others say fine-tuning works better: Margatina et al. [92] propose AL with Pretrained LMs: Starting with a pretrained LM $\mathcal{P}(x;W_{0})$, the model undergoes task-adaptive pretraining on the unlabeled pool $D_{\text{pool}}$ to obtain $\mathcal{P}_{\text{TAPT}{}}(x;W_{0}^{\prime})$. In each AL round, a classifier $\mathcal{M}_{i}$ is initialized from $\mathcal{P}_{\text{TAPT}{}}$ and fine-tuned on the labeled set $D_{\text{lab}}$. A batch is selected using the acquisition function $a$, and the process repeats.
Active query k-means offers another option beyond traditional methods [94]. Many sampling strategies are proposed, such as uncertainty sampling, query-by-committee, and Penalized Min-Max-selection [5, 93, 94].
AL in NLP still has open problems. One issue is the lack of strong comparisons between methods. This makes it hard to know which strategy is best [95]. Also, labeling NLP data takes time and money, especially for large tasks [5, 93]. Adapting language models to new tasks is also not easy and needs more research [92]. Recent studies look at active learning for sequence labeling [97] and machine translation [98]. These show that active learning can help in many NLP tasks.
To sum up, active learning shows strong potential in NLP. It helps save labels and improve results. But more research is needed to solve current challenges. New methods like active query k-means and better model adaptation can lead to better results [5, 92, 93]. In the future, active learning could help build more accurate and cost-efficient NLP models [95, 94].
Algorithm 5 Active Learning with Pretrained LMs [92]
@@
Figure 6: Illustration of AADA [99] algorithm. The algorithm operates within the unsupervised domain adaptation framework. Initially, the authors train a model using the labeled source data ($L_{s}$) and the unlabeled target data ($U_{t}$), leveraging a domain adversarial loss. In each subsequent iteration, the authors select informative samples from the target domain based on importance weighting and acquire their labels. The model is then re-trained using the combined labeled data ($L_{s}\cup L_{t}$) along with the remaining unlabeled target data ($U_{t}$).
@@
Figure 7: An overview of the interactive framework by Lee et al. [100]. Unlike prior approaches that focus solely on optimizing the system objective (blue modules), their method also incorporates the user objective (green line), enabling a joint optimization of these potentially conflicting goals (yellow block).
@@
Domain Adaptation and Transfer Learning in Active Learning: domain adaptation and transfer learning enable models to perform competently in different domains and tasks. Several recent works have attempted to resolve challenges such as domain shift and knowledge transfer. Active Adversarial Domain Adaptation (AADA) [99] utilizes adversarial domain alignment and importance sampling within a unified framework. It chooses samples based on their uncertainty and diversity, indicating that these two aspects play a crucial role in active learning.
@@
The domain adaptation survey for visual applications [101] provides a decent summary of techniques in image classification and object detection. It covers both shallow and deep learning methods. There exists a survey on the theory of domain adaptation [102] that deals with learning bounds and theoretical guarantees. It indicates that there should be more analysis and theory for the field.
@@
A number of new approaches have also been proposed. Divide and Adapt (DiaNA) [103] constructs four target groups of different transferability. It adapts each of these groups using a particular approach, and the performance improves. Adaptive Feature Norm [104] demonstrates that features with higher norms transfer optimally. This approach outperforms other state-of-the-art methods for unsupervised domain adaptation.
@@
Most of the papers concur that domain adaptation plays a crucial role in active learning. They also mention how one should select quality samples for labeling. Transfer learning helps transfer knowledge across tasks and domains. Uncertainty and diversity are shared techniques for sampling. We observe the same in AADA [99], DiaNA [103], and other such tasks like Discriminative Active Learning [105] and Domain Adversarial Reinforcement Learning [106].
@@
Varying technical methods are utilized. Adversarial alignment [99] is applied by AADA. Importance sampling and Gaussian Mixture Models are utilized by DiaNA [103]. Adaptive Feature Norm adjusts the norms of features so that domains become similar [104]. Learning to Adapt [107] and Domain-Invariant Feature Learning [108] are other methods that reduce domain shift. Despite progress, there remain challenges. One of them is negative transfer, in which performance degrades following adaptation [101, 102].
@@
Challenges remain in domain adaptation: Certain methods are slow or do not generalize effectively to big data [104, 103]. The frequent lack of theoretical analysis constrains our knowledge [102]. To treat overfitting or underfitting, we require regularization and thoughtful model selection [99, 103].
@@
In conclusion, transfer learning and domain adaptation are central to enabling active learning in different tasks. Several effective techniques have long been proposed for addressing domain shift and knowledge transfer. There remain some problems that still require work. Fresh ideas, such as in Towards Domain Adaptation [109] and the Survey on Transfer Learning [110], will keep advancing the field. All these advancements are crucial for making active learning more effective and useful.
@@
Figure 8: Illustration of the DEAL [111]: difficulty-aware active learning framework for semantic segmentation. The architecture consists of two main branches: the first follows a standard semantic segmentation network, while the second incorporates a probability attention module followed by a $1\times 1$ convolution. Here, $D^{a}$ and $D^{u}$ represent the labeled and unlabeled datasets, respectively, and $D^{s}$ denotes a selected subset from $D^{u}$. Probability maps before and after the attention mechanism are denoted by $P$ and $Q$, respectively. The framework optimizes two loss functions, $\mathcal{L}_{\text{seg}}$ and $\mathcal{L}_{\text{dif}}$ which also introduced in this work [111].
@@
Real-World Applications and Case Studies of Active Learning: in recent years, many papers have studied how active learning works in real-world applications. A simulation study in [1] tests different active learning methods in various situations. It shows that we need clear and accurate metrics to measure performance. Another paper [4] introduces actionable metrics. These help people in industry better understand and evaluate active learning, making it easier to trust and apply. In education, [100] shows how active learning can improve student engagement. These examples show that active learning is important for building strong AI models in different areas.
@@
Balancing system performance with user needs is also an important topic. For example, [100] proposes a method that considers both model quality and user demands. It works well in an educational setting. This kind of user-focused approach helps make active learning more reliable in real-world use. But other papers focus more on improving system performance, showing a trade-off between system goals and user experience [3]. Simulation experiments like [112] help test model performance. Query strategies from [5] are key to selecting useful samples.
@@
However, there exist many challenges. One big challenge is exploration—active learning sometimes struggles to search the full data space [113]. Also, picking the best samples to label is not easy [15]. Some papers discuss how to handle AutoML systems [3], and how to bring research ideas into practical industry use [114].
@@
Active learning can be used in many industries. It also connects with other areas like reinforcement learning [11] and human-like learning [15]. As the research grows, we must solve open problems to make active learning useful. We need better metrics, stronger technical solutions, and a good balance between model performance and user needs. Future progress depends on solving these challenges [9, 39].
@@
Active Learning for Specialized Tasks: AL is used more often in specialized tasks like semantic segmentation, object detection, and recognition nowadays. The goal is to get good results using only a small amount of labeled data. Many recent studies say that we need better and more realistic ways to evaluate active learning methods [41], [58]. For example, the research shows that deep active learning may not always perform better than random sampling, and results can change a lot depending on how the model is trained [41]. Another work, ALBench, builds a benchmark to help researchers compare methods fairly and reproduce results in object detection [58].
@@
In semantic segmentation, active learning tries to focus on areas that are harder to segment. This helps improve results in those difficult regions [111]. The DEAL method [111] learns how to estimate which areas are hard to segment using pixel-wise attention scores. This shows the importance of designing active learning methods for the specific task. Other papers combine active learning with semi-supervised learning to gain performance on segmentation tasks [115], [116]. But the performance depends on the task and the dataset.
@@
In object detection, active learning is used to lower the cost of labeling and improve model accuracy [117], [118]. The ALWOD [118] shows a method for active learning in weakly supervised object detection. It proves that active learning can help in this area too. Another study looks at how well low-cost methods work for object detection [119]. It points out the need for more efficient strategies.
@@
Several papers also talk about real-world problems. They say active learning must adapt to changes in data and task conditions [114], [120]. Discwise AL for LiDAR semantic segmentation [121] shows how active learning can be used in more types of data. The Best Practices paper gives advice on how to choose the right active learning method [120]. It considers factors like the type of data, whether semi-supervised learning is used, and how many labels are available.
@@
To summarize, active learning is now widely used in tasks like segmentation, detection, and recognition. It helps improve results when labeled data is limited. But we still need better benchmarks [58], [41], more realistic tests, and smarter methods that handle specific tasks and real-world problems. Combining active learning with semi-supervised learning is one promising direction. Still, more research is needed to fix current issues and build faster, more reliable active learning systems [41], [58].

### v2

@@ IV Evaluations
Evaluation Metrics and Benchmarks for Active Learning: there have been calls from recent research for sound and reproducible evaluation, with appropriate re-benchmarking of available benchmarks in order to yield more sound results [2306.08954]. New benchmarks like ALdataset [2010.08161], OpenAL [2304.05246], and CDALBench [2408.00426] facilitate end-to-end comparisons.
A central theme is the recommendation of practical metrics for restoring trust in AL among such studies [abraham2020rebuilding]. Metrics like interpretability metrics and statistical analysis techniques provide greater insight into AL performance. Cross-domain validation and large-scale reproducibility guarantee that methods generalize from one dataset [2408.00426]. Furthermore, interpretability and transparency are essential. Studies demand metrics that explain model behavior [2304.05246, abraham2020rebuilding]. Open experimentation environments and shared code enhance reproducibility [2306.08954, 2010.08161].
Perspectives on metrics and settings differ. Some researchers aim to design new benchmarks and evaluation metrics [2010.08161], e.g., OpenAL, proposed by Jonas et al. [2304.05246]. Others advocate using existing metrics with new analysis techniques [abraham2020rebuilding]. Experiment settings also vary between studies, from differences in the choice of datasets, oracle implementation, and repetition count [2408.00426, 2306.08954]. Methodologies vary from benchmark design to statistical analysis. The design of new benchmarks like ALdataset and OpenAL necessitates attention to experimental settings and evaluation metrics [2010.08161, 2304.05246]. Open-sourced implementations as well as efficient oracle implementations are also important in enabling the comparison of AL approaches [2306.08954, 2408.00426].
In spite of the advancements made in the construction of evaluation metrics and benchmarks for AL, there are still limitations and challenges. A deficiency in the standardization of benchmarks and evaluation metrics prevents comparison of AL approaches, and variability in the experimental conditions can give rise to conflicting results [2306.08954, 2010.08161, 1108.0453, 1504.02406]. The necessity of performing large-scale evaluation with repeated runs highlights the importance of making thorough and trustworthy assessments of AL approaches [2408.00426].
In summary, working towards evaluation metrics and benchmarks for AL remains an important research focus, with themes such as the necessity for trustworthy benchmarks, interpretability, cross-domain evaluation, and reproducibility. Metrics and setups for experiments vary among researchers. The technical approaches from these works provide a strong base for subsequent research. Overcoming the limitations and challenges of these works will be vital in furthering our knowledge of AL strategies and their applications in other domains [grimm2015self, iyyer2017amazing].
@@ V Applications
Applications of Active Learning in Computer Vision: AL is becoming more popular in computer vision because it helps reduce the cost of data labeling [takezoe2023deep]. Many recent studies focus on lowering annotation costs while keeping training efficient [tuia2012learning, 1908.02454]. For example, learning how confident a user is when labeling can help avoid asking for labels that are not necessary, which improves the overall process [tuia2012learning]. Some methods first use weak labels and only request strong labels later. This helps save effort in tasks like object detection [1908.02454].
@@
Active learning has many applications in computer vision, e.g. image classification, object detection, and semantic segmentation [2012.04225, 2207.13339]. New methods aim to make active learning more effective and faster, including uncertainty sampling, query-by-committee, and reinforcement learning [2108.05595, kyrkou2020imitation]. Several papers also discuss how to evaluate active learning better. They suggest new evaluation metrics and testing protocols for realistic comparisons [10.1007/978-3-642-41398-8_16, 1912.05361].
@@
Take Greedy Active Learning (GAL) for example. Bar et al. [2412.02310] propose a greedy-based active learning framework that operates in two stages during each AL cycle. Denote the classifier at iteration $t$ as $\mathcal{C}_{t}$. In the initial stage, a candidate subset $X_{c}\subseteq X_{u}$ of size $K:=|X_{c}|$ is selected from the unlabeled data pool. This subset may either encompass the entire unlabeled set or be filtered based on the top-K relevance scores. Due to inherent class imbalance, $X_{c}$ typically contains many irrelevant samples. In the following stage, a batch $X_{b}\subset X_{c}$ is chosen using an AL strategy. A human oracle then labels the instances in $X_{b}$, updating the labeled set $(X_{l},\mathcal{Y}_{l})$ with their features and labels. A new classifier $\mathcal{C}_{t+1}$ is trained on this augmented dataset for the next iteration.
@@
Researchers are questioning traditional active learning methods. Some propose new frameworks that combine semi-supervised learning and data augmentation [1912.05361]. Others explore combining active learning with transfer learning and self-supervised learning [2301.01531, qu2025recent]. Deep learning models like CNNs and visual transformers are also widely used in active learning for computer vision tasks [jeon2017active, 2106.03801].
@@
Despite recent progress, some challenges still exist. It is hard to scale active learning to very large datasets or complex problems [parvaneh2022active, 2412.02310]. Also, poor-quality data or labels can hurt performance. Researchers are working on more efficient algorithms to solve these issues [2202.01402]. High-quality datasets and better annotation protocols are also important for success [bosser2021model].
@@
In conclusion, AL plays an important role in computer vision. It helps reduce labeling costs and improve learning efficiency. Recent work explores new evaluation methods and combines active learning with other learning techniques. Researchers continue to improve active learning approaches [bang2024active, 2406.03845] to mitigate the challenges. The field is growing, and active learning will likely become even more important in future computer vision systems.
Algorithm 3 Greedy Active Learning (GAL) Algorithm [2412.02310]
@@
   $x^{*},\hat{l}^{*}\leftarrow\textsc{Next}(\mathcal{X}_{c},\mathcal{X}_{l},\mathcal{Y}_{l})$ $\triangleright$ Find the point that maximizes impact $\mathcal{S}$
@@
   $\mathcal{S}_{i},\hat{l}_{i}\leftarrow\psi(x_{i},\mathcal{X}_{l},\mathcal{Y}_{l})$ $\triangleright$ Acquisition function (see Alg. 4)
@@
  return $x_{i^{*}},\hat{l}_{i^{*}}$ $\triangleright$ Return optimal point and pseudo label
@@
Algorithm 4 Acquisition Functions [2412.02310]
[…]
@@
[…]
@@
Active Learning for Natural Language Processing Tasks: many papers show AL can help improve data efficiency and model accuracy [zhang2022survey, 2104.08320, 1810.03450]. One important idea is adapting pre-trained language models to the specific task. Training more on unlabeled data can improve results and reduce the need for labels [2104.08320]. For example, ensemble methods like Majority-CRF [1810.03450] are used in natural language understanding. They reduce error rates by 6.6%–9% compared to random sampling. Another method, active query k-means, also helps increase accuracy and cut training costs in text classification [jiang2021application].
Model adaptation is a crucial topic in NLP active learning. It shows the importance of learning strategies for pre-trained models [zhang2022survey, 2104.08320]. Ensemble methods are popular too, as they help pick the most useful samples [1810.03450]. Semi-supervised learning is often used with active learning to reduce labeling needs [kohl2024scoping, jiang2021application]. Beatty et al. show that choosing when to stop active learning using labeled vs. unlabeled data is important, especially for text classification [beatty2019use]. Good evaluation tools, e.g., datasets and metrics, are also important. Most papers use F1-score to measure performance [kohl2024scoping]. The corresponding equations in pseudocodes are:
@@
There are different opinions on which query strategy is best. Some focus on exploitation, some on exploration, and others mix both [kohl2024scoping]. Some researchers prefer continuous training on unlabeled data to adapt Language Models (LMs). Others say fine-tuning works better: Margatina et al. [2104.08320] propose AL with Pretrained LMs: Starting with a pretrained LM $\mathcal{P}(x;W_{0})$, the model undergoes task-adaptive pretraining on the unlabeled pool $D_{\text{pool}}$ to obtain $\mathcal{P}_{\text{TAPT}{}}(x;W_{0}^{\prime})$. In each AL round, a classifier $\mathcal{M}_{i}$ is initialized from $\mathcal{P}_{\text{TAPT}{}}$ and fine-tuned on the labeled set $D_{\text{lab}}$. A batch is selected using the acquisition function $a$, and the process repeats.
Active query k-means offers another option beyond traditional methods [jiang2021application]. Many sampling strategies are proposed, such as uncertainty sampling, query-by-committee, and Penalized Min-Max-selection [zhang2022survey, 1810.03450, jiang2021application].
AL in NLP still has open problems. One issue is the lack of strong comparisons between methods. This makes it hard to know which strategy is best [kohl2024scoping]. Also, labeling NLP data takes time and money, especially for large tasks [zhang2022survey, 1810.03450]. Adapting language models to new tasks is also not easy and needs more research [2104.08320]. Recent studies look at active learning for sequence labeling [humeniuk2023ambiegen] and machine translation [virtanen2022product]. These show that active learning can help in many NLP tasks.
To sum up, active learning shows strong potential in NLP. It helps save labels and improve results. But more research is needed to solve current challenges. New methods like active query k-means and better model adaptation can lead to better results [zhang2022survey, 2104.08320, 1810.03450]. In the future, active learning could help build more accurate and cost-efficient NLP models [kohl2024scoping, jiang2021application].
Algorithm 5 Active Learning with Pretrained LMs [2104.08320]
@@
Figure 6: Illustration of AADA [su2020active] algorithm. The algorithm operates within the unsupervised domain adaptation framework. Initially, the authors train a model using the labeled source data ($L_{s}$) and the unlabeled target data ($U_{t}$), leveraging a domain adversarial loss. In each subsequent iteration, the authors select informative samples from the target domain based on importance weighting and acquire their labels. The model is then re-trained using the combined labeled data ($L_{s}\cup L_{t}$) along with the remaining unlabeled target data ($U_{t}$).
@@
Figure 7: An overview of the interactive framework by Lee et al. [2005.04470]. Unlike prior approaches that focus solely on optimizing the system objective (blue modules), their method also incorporates the user objective (green line), enabling a joint optimization of these potentially conflicting goals (yellow block).
@@
Domain Adaptation and Transfer Learning in Active Learning: domain adaptation and transfer learning enable models to perform competently in different domains and tasks. Several recent works have attempted to resolve challenges such as domain shift and knowledge transfer. Active Adversarial Domain Adaptation (AADA) [su2020active] utilizes adversarial domain alignment and importance sampling within a unified framework. It chooses samples based on their uncertainty and diversity, indicating that these two aspects play a crucial role in active learning.
@@
The domain adaptation survey for visual applications [1702.05374] provides a decent summary of techniques in image classification and object detection. It covers both shallow and deep learning methods. There exists a survey on the theory of domain adaptation [2004.11829] that deals with learning bounds and theoretical guarantees. It indicates that there should be more analysis and theory for the field.
@@
A number of new approaches have also been proposed. Divide and Adapt (DiaNA) [huang2023divide] constructs four target groups of different transferability. It adapts each of these groups using a particular approach, and the performance improves. Adaptive Feature Norm [xu2019larger] demonstrates that features with higher norms transfer optimally. This approach outperforms other state-of-the-art methods for unsupervised domain adaptation.
@@
Most of the papers concur that domain adaptation plays a crucial role in active learning. They also mention how one should select quality samples for labeling. Transfer learning helps transfer knowledge across tasks and domains. Uncertainty and diversity are shared techniques for sampling. We observe the same in AADA [su2020active], DiaNA [huang2023divide], and other such tasks like Discriminative Active Learning [zhou2021discriminative] and Domain Adversarial Reinforcement Learning [chen2020domain].
@@
Varying technical methods are utilized. Adversarial alignment [su2020active] is applied by AADA. Importance sampling and Gaussian Mixture Models are utilized by DiaNA [huang2023divide]. Adaptive Feature Norm adjusts the norms of features so that domains become similar [xu2019larger]. Learning to Adapt [dakic2017macroscopic] and Domain-Invariant Feature Learning [1310.5195] are other methods that reduce domain shift. Despite progress, there remain challenges. One of them is negative transfer, in which performance degrades following adaptation [1702.05374, 2004.11829].
@@
Challenges remain in domain adaptation: Certain methods are slow or do not generalize effectively to big data [xu2019larger, huang2023divide]. The frequent lack of theoretical analysis constrains our knowledge [2004.11829]. To treat overfitting or underfitting, we require regularization and thoughtful model selection [su2020active, huang2023divide].
@@
In conclusion, transfer learning and domain adaptation are central to enabling active learning in different tasks. Several effective techniques have long been proposed for addressing domain shift and knowledge transfer. There remain some problems that still require work. Fresh ideas, such as in Towards Domain Adaptation [baldassi2016unreasonable] and the Survey on Transfer Learning [tan2018survey], will keep advancing the field. All these advancements are crucial for making active learning more effective and useful.
@@
Figure 8: Illustration of the DEAL [xie2020deal]: difficulty-aware active learning framework for semantic segmentation. The architecture consists of two main branches: the first follows a standard semantic segmentation network, while the second incorporates a probability attention module followed by a $1\times 1$ convolution. Here, $D^{a}$ and $D^{u}$ represent the labeled and unlabeled datasets, respectively, and $D^{s}$ denotes a selected subset from $D^{u}$. Probability maps before and after the attention mechanism are denoted by $P$ and $Q$, respectively. The framework optimizes two loss functions, $\mathcal{L}_{\text{seg}}$ and $\mathcal{L}_{\text{dif}}$ which also introduced in this work [xie2020deal].
@@
Real-World Applications and Case Studies of Active Learning: in recent years, many papers have studied how active learning works in real-world applications. A simulation study in [10.1007/978-3-642-41398-8_16] tests different active learning methods in various situations. It shows that we need clear and accurate metrics to measure performance. Another paper [abraham2020rebuilding] introduces actionable metrics. These help people in industry better understand and evaluate active learning, making it easier to trust and apply. In education, [2005.04470] shows how active learning can improve student engagement. These examples show that active learning is important for building strong AI models in different areas.
@@
Balancing system performance with user needs is also an important topic. For example, [2005.04470] proposes a method that considers both model quality and user demands. It works well in an educational setting. This kind of user-focused approach helps make active learning more reliable in real-world use. But other papers focus more on improving system performance, showing a trade-off between system goals and user experience [bosser2021model]. Simulation experiments like [1408.1319] help test model performance. Query strategies from [zhang2022survey] are key to selecting useful samples.
@@
However, there exist many challenges. One big challenge is exploration—active learning sometimes struggles to search the full data space [1206.5274]. Also, picking the best samples to label is not easy [2011.03733]. Some papers discuss how to handle AutoML systems [bosser2021model], and how to bring research ideas into practical industry use [1907.00038].
@@
Active learning can be used in many industries. It also connects with other areas like reinforcement learning [2201.03947] and human-like learning [2011.03733]. As the research grows, we must solve open problems to make active learning useful. We need better metrics, stronger technical solutions, and a good balance between model performance and user needs. Future progress depends on solving these challenges [2306.08001, 2403.14800].
@@
Active Learning for Specialized Tasks: AL is used more often in specialized tasks like semantic segmentation, object detection, and recognition nowadays. The goal is to get good results using only a small amount of labeled data. Many recent studies say that we need better and more realistic ways to evaluate active learning methods [1912.05361], [2207.13339]. For example, the research shows that deep active learning may not always perform better than random sampling, and results can change a lot depending on how the model is trained [1912.05361]. Another work, ALBench, builds a benchmark to help researchers compare methods fairly and reproduce results in object detection [2207.13339].
@@
In semantic segmentation, active learning tries to focus on areas that are harder to segment. This helps improve results in those difficult regions [xie2020deal]. The DEAL method [xie2020deal] learns how to estimate which areas are hard to segment using pixel-wise attention scores. This shows the importance of designing active learning methods for the specific task. Other papers combine active learning with semi-supervised learning to gain performance on segmentation tasks [2210.08403], [2203.10730]. But the performance depends on the task and the dataset.
@@
In object detection, active learning is used to lower the cost of labeling and improve model accuracy [goupilleau2021active], [wang2023alwod]. The ALWOD [wang2023alwod] shows a method for active learning in weakly supervised object detection. It proves that active learning can help in this area too. Another study looks at how well low-cost methods work for object detection [probst2022evaluating]. It points out the need for more efficient strategies.
@@
Several papers also talk about real-world problems. They say active learning must adapt to changes in data and task conditions [1907.00038], [mittal2023best]. Discwise AL for LiDAR semantic segmentation [unal2023discwise] shows how active learning can be used in more types of data. The Best Practices paper gives advice on how to choose the right active learning method [mittal2023best]. It considers factors like the type of data, whether semi-supervised learning is used, and how many labels are available.
@@
To summarize, active learning is now widely used in tasks like segmentation, detection, and recognition. It helps improve results when labeled data is limited. But we still need better benchmarks [2207.13339], [1912.05361], more realistic tests, and smarter methods that handle specific tasks and real-world problems. Combining active learning with semi-supervised learning is one promising direction. Still, more research is needed to fix current issues and build faster, more reliable active learning systems [1912.05361], [2207.13339].